40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Message ID: 366     Entry time: Mon Mar 10 02:05:08 2008
Author: rob 
Type: Update 
Category: Locking 
Subject: DRMI+2ARMs working better 

Some encouraging progress on the locking front tonight. After the work on the DRM loops last week and a review of the settings for initial lock acquisition (loop gains, tickle amplitude, filter states, so on), the DRMI+2ARMS locking is working pretty well. That's to say, it takes from 5-15 minutes generally for the IFO to lock in the offset CARM state, with the arm powers at 0.5. It's then possible to raise the arm powers slightly, and handing off control of CARM to MCL works at low power, but engaging the AO path (using PO_DC as an error signal) is not working so well. Taking swept sines indicates that the PO_DC should be a good error signal. The next good thing to try might be just using PO_DC as an error signal for the length path, without using the AO path at all, to see if it's something in the hardware.
ELOG V3.1.3-